4y2+9=0
16+1y+9=0
y=-25
Is this correct? Thank you

Answer by Nate(3500)

4y^2 + 9 = 0
4y^2 = -9
y^2 = -9/4
y = +- sqrt(-9/4)
y = +- 3i/2